# Insights for Early Dark Energy with Big Bang Nucleosynthesis

### v1

@@ I Introduction
Big Bang Nucleosynthesis (BBN) is a cornerstone of modern cosmology. As the earliest process in the universe that can currently be tested against direct observations, BBN provides one of our most powerful windows into the physics of the early universe. There exists a well-established theoretical framework connecting the light element abundances to fundamental cosmological parameters; see Refs. Cyburt:2015mya; Fields:2019pfx; Grohs:2023voo; ParticleDataGroup:2024cfk for recent reviews. The predictions of standard BBN for the primordial abundances of the light elements, particularly helium-4 and deuterium, are in remarkable agreement with measurements, confirming the robustness of the Standard Model of particle physics and cosmology ParticleDataGroup:2024cfk. At the same time, because the abundance of these elements depends sensitively on the strength of fundamental interactions and on the cosmic history, BBN offers a unique probe of new physics in the early universe.
Recent years have produced rapid progress in both observational and theoretical aspects of BBN. Measurements of the primordial abundances of helium-4 Aver:2020fon; Valerdi:2019beb; Fernandez:2019hds; Kurichin:2021ppm; Hsyu:2020uqb; Valerdi:2021abc; Aver:2021rwi; ParticleDataGroup:2024cfk and deuterium Cooke:2013cba; Cooke:2016rky; Riemer-Sorensen:2014aoa; Riemer-Sorensen:2017pey; Balashev:2015hoe; Zavarygin:2017cov; Cooke:2017cwo; ParticleDataGroup:2024cfk have reached unprecedented precision, yielding abundances consistent with Standard Model predictions at the percent level. Experimental efforts, such as those by the LUNA collaboration Mossa:2020gjc, have reduced uncertainties in the nuclear reaction rates that govern BBN yields. Moreover, observations of the cosmic microwave background (CMB) now tightly constrain the baryon density Planck:2018vyg; AtacamaCosmologyTelescope:2025blo; SPT-3G:2025bzu, eliminating one of the largest sources of uncertainty in comparing BBN theory with data. Numerical computation of the predictions of BBN for a range of cosmological scenarios can be achieved through various publicly available software packages Arbey:2011nf; Arbey:2018zfh; Pisanti:2007hk; Consiglio:2017pot; Pitrou:2019nub; Giovanetti:2024zce. Together, these advances establish BBN as an incisive tool for testing physics beyond the Standard Model.
@@
Several avenues have been explored in this direction Allahverdi:2020bys. For example, small-scale structure places constraints on modifications to the expansion history at early times Erickcek:2011us; Delos:2018ueo, while gravitational wave backgrounds Caprini:2018mtu are, in principle, sensitive to the equation of state during very early epochs Watanabe:2006qe; Barenboim:2016mjm. Studies of the effects of the pre-recombination expansion history using principal component methods (Samsing:2012qx) and constraints from the matter power spectrum on early dark energy (Sobotka:2024ixo) have provided complementary insights into possible deviations from the standard expansion history. Yet BBN offers a particularly powerful and direct probe, since light element abundance yields are directly sensitive to the expansion rate during the first few minutes after the Big Bang.
Previous studies have shown that BBN abundances can be used to constrain the effective equation of state during nucleosynthesis Carroll:2001bv; Arbey:2009gt, but the steadily improving precision of abundance measurements now makes it possible to go further. Recently, Ref. (An:2023buh) explored a model-independent reconstruction of the expansion and thermal histories constrained by the light element abundances. In this paper, we use the latest abundance data to derive general constraints on the expansion history during BBN. However, unlike Ref. (An:2023buh), we employ principal component analysis to extract the most robust, model-independent features of these constraints, allowing us to place limits on broad classes of early-universe scenarios, including those motivated by early dark energy. In addition, we revisit the long-standing lithium problem, the persistent discrepancy between predicted and observed primordial lithium abundances Fields:2011zzb; ParticleDataGroup:2024cfk. While it is well established that uncertainties in nuclear reaction rates cannot resolve this discrepancy ParticleDataGroup:2024cfk, we investigate whether a general modification to the expansion history could mitigate the tension.
@@ II Big Bang Nucleosynthesis
As such, BBN plays a pivotal role in constraining both standard cosmological parameters and potential new physics beyond the Standard Model, such as decaying particles, asymmetric reheating, or varying fundamental constants ParticleDataGroup:2024cfk. Among the most robust predictions are those for deuterium and helium-4, whose abundances can now be measured with high precision in nearly pristine astrophysical environments. The concordance between BBN predictions and observational data for these elements offers strong support for the standard cosmological model. However, the observed primordial abundance of lithium-7 remains significantly lower than theoretical expectations, a discrepancy known as the “lithium problem,” which may point to underestimated observational errors or hints of new physics Fields:2011zzb; Fields:2019pfx; ParticleDataGroup:2024cfk.
@@ II.1 Observational Status
Recent observations of primordial light element abundances continue to support the standard cosmological model for most isotopes while highlighting persistent anomalies. We use the Particle Data Group (PDG) ParticleDataGroup:2024cfk recommended values of observed abundance ratios throughout this work, which are summarized in Table 1. Deuterium remains the most accurately measured primordial isotope, with high-redshift quasar absorption systems yielding D/H = $(2.547\pm 0.029)\times 10^{-5}$, in excellent agreement with BBN predictions when using the CMB-derived baryon density. Similarly, helium-4 determinations from extragalactic HII regions, aided by improved emissivity corrections and infrared line data, support a mass fraction $Y_{p}=0.245\pm 0.003$, matching theoretical expectations. The primordial abundance of helium-3, however, remains poorly constrained due to its detection being limited to metal-rich Galactic environments. Lithium-6, once thought to be overabundant, is now strongly constrained in both Galactic halo stars and the Small Magellanic Cloud (SMC), with no confirmed detections, effectively ruling out a primordial 6Li plateau and supporting the case for stellar depletion.
Lithium-7 remains a significant challenge to the standard BBN paradigm. The Spite plateau in metal-poor halo stars yields a 7Li/H ratio of $(1.6\pm 0.3)\times 10^{-10}$, a factor of three below the BBN+CMB prediction of $(4.72\pm 0.7)\times 10^{-10}$. New observations reveal dispersion and decreasing 7Li levels at metallicities [Fe/H] $\lesssim-2.7$, pointing to stellar depletion as a plausible explanation Fields:2022mpw. In parallel, stringent upper limits on 6Li/7Li, such as <0.007<0.007 in HD 84937 and <0.1<0.1 in the SMC’s Sk 143 sightline, suggest significant 6Li depletion. Since 6Li is more fragile than 7Li, these findings imply that comparable 7Li destruction is astrophysically feasible, potentially resolving the discrepancy without requiring new physics.
Yet, extragalactic evidence challenges this resolution. The detection of gas-phase 7Li in the SMC with an abundance $A(^{7}\mathrm{Li})\approx 2.2$, unaffected by stellar processes, still falls below BBN expectations Molaro:2024wxa. This implies a primordial shortfall that stellar depletion alone cannot explain. While Galactic 6Li non-detections support an astrophysical solution, the SMC results suggest a possible need for physics beyond the Standard Model. Although stellar depletion offers a compelling explanation within the Milky Way, the lithium problem may ultimately point to deeper cosmological questions.
@@
Table 1: Primordial light element abundances: observations vs. Standard BBN predictions (assuming $\eta_{10}=6.143\pm 0.190$). This work uses the PDG ParticleDataGroup:2024cfk recommended values of observed abundance ratios.
@@ III.1 BBN Calculations
To study the effects of a non-standard expansion history on BBN, we use a modified version of AlterBBN (Arbey:2011nf; Arbey:2018zfh). This publicly available code numerically solves the coupled differential equations governing the evolution of light nuclei in the early universe. Boltzmann solvers like AlterBBN are essential tools in cosmology because they compute the time evolution of particle species based on their interactions, decay rates, and the universe’s expansion history. Specifically for BBN, AlterBBN integrates the nuclear reaction network for light elements, such as deuterium, helium-3, helium-4, and lithium-7, while simultaneously solving the Friedmann equation that determines the expansion rate as a function of time $H(t)$.
@@ III.3 Fisher Matrix and Principal Component Analysis
These principal components provide a powerful basis for model-independent constraints: any arbitrary EDE history can be projected onto the eigenbasis, allowing one to reconstruct the constrained and unconstrained features of the model in terms of measurable quantities. The plot of the three eigenmodes are shown in Figure 3. The most strongly constrained mode is degenerate with a uniform increase to the radiation density. However, as seen in the second and third eigenmodes, the primordial abundances are also sensitive to temperature-dependent changes in the expansion history, especially those that affect expansion in the temperature range $10^{9}~\mathrm{K}>T_{\gamma}>3\times 10^{8}$ K. Qualitatively similar conclusions were drawn from the binned expansion history analysis presented in Ref. An:2023buh, where the authors also found that modifications to the expansion rate around this temperature range have the most significant impact on light element abundances. The agreement between the PCA-based and binned analyses provides a strong validation that the constraints identified here are robust to the choice of parametrization.
@@ III.4 Markov Chain Monte Carlo Analysis
To explore the full posterior distribution of EDE histories and relevant cosmological parameters, we perform a Markov Chain Monte Carlo (MCMC) analysis using the Cobaya framework Torrado:2020dgo. Cobaya (Code for Bayesian Analysis) is a modular, extensible Python-based platform for efficient Bayesian inference, particularly well suited to hierarchical physical models with complex interdependencies.
@@
  • $\eta_{10}\equiv 10^{10}\times\frac{n_{b}}{n_{\gamma}}$: the baryon-to-photon ratio, with a Gaussian prior $\eta_{10}=6.104\pm 0.058$, consistent with the constraint derived from CMB observations Planck:2018vyg; ParticleDataGroup:2024cfk,
  • $\tau_{n}$: the neutron lifetime, with a Gaussian prior $\tau_{n}=878.4\pm 0.5$ s, consistent with laboratory measurements ParticleDataGroup:2024cfk,
@@
Each set of parameters defines a complete model of early dark energy. Note that the first principal component of the dark energy density is degenerate with that of the radiation density, typically parameterized by $N_{\mathrm{eff}}$, so we did not separately vary $N_{\mathrm{eff}}$ in our analysis. Several previous studies have shown how observations of primordial abundances can be used to constrain the light relic density ParticleDataGroup:2024cfk; Cyburt:2015mya; Fields:2019pfx; Lague:2019yvs; Dvorkin:2022jyg; Yeh:2022heq.
@@
Sampling is performed using Cobaya’s built-in Metropolis-Hastings algorithm, which automatically groups parameters into fast and slow blocks based on the computational cost of the components that depend on them Lewis:2002ah; Lewis:2013hha; Neal:2005uqf; Torrado:2020dgo. This automatic parameter blocking and caching of intermediate results greatly improves sampling efficiency. Chain convergence is monitored using the Gelman-Rubin R1R-1 statistic, with a typical stopping criterion of R1<0.01R-1<0.01.
Posterior samples are processed using GetDist Lewis:2019xzd, which interfaces seamlessly with Cobaya’s output, providing marginal distributions, credible intervals, and covariance matrices. In particular, we analyze the posterior distributions of the eigenmode amplitudes $\alpha_{i}$, and project the results back into temperature space to reconstruct allowed EDE histories and identify epochs most sensitive to BBN constraints. We use the observational constraints on primordial abundances from the PDG ParticleDataGroup:2024cfk, summarized above in Table 1.
@@ IV Results
Figure 4 shows the allowed EDE histories consistent with the observed primordial helium-4 and deuterium abundances. In this scenario, the EDE component is generally constrained to remain subdominant throughout the epoch of BBN, except in the narrow temperature range $3\times 10^{8}\,\mathrm{K}\lesssim T\lesssim 5\times 10^{8}\,\mathrm{K}$, where a brief period of dominance is permitted. A modest secondary enhancement is also visible around or just after the deuterium bottleneck, indicating a localized allowance for additional EDE energy density during that stage. These results are broadly consistent with the binned reconstruction of the expansion history presented in Ref. An:2023buh, confirming that the light-element abundances primarily constrain deviations in the Hubble rate near the onset of nucleosynthesis. The PCA framework developed here extends that earlier analysis by providing a continuous, model-independent mapping of the allowed expansion histories, revealing not only where the expansion rate is tightly constrained but also which temperature ranges remain weakly bounded by current data.
@@
The analysis presented here also allows us to determine how future data might impact our ability to constrain the early expansion history. Figure 5 presents the corresponding histories obtained after including a forecasted $1\%$ constraint on the helium-3 abundance. The addition of this prior produces minimal qualitative change in the allowed EDE behavior. As before, EDE remains largely subdominant, with the same narrow window of possible dominance between $3\times 10^{8}$ K and $5\times 10^{8}$ K, and a slight spike near or following the deuterium bottleneck.
In contrast, Figure 6 illustrates the allowed EDE histories when only the helium mass fraction and deuterium abundance are constrained. At the same time, the prior on $\eta_{10}$ is relaxed to a flat range $[1,20]$. In this case, EDE is permitted to dominate throughout the entire BBN epoch, reaching energy densities several orders of magnitude higher than in the previous scenarios. However, this freedom is accompanied by extreme values of $\eta_{10}$ that are in strong tension with the precise constraints from CMB observations, highlighting the importance of an informed prior on the baryon-to-photon ratio.

### v2

@@ I Introduction
Big Bang Nucleosynthesis (BBN) is a cornerstone of modern cosmology. As the earliest process in the universe that can currently be tested against direct observations, BBN provides one of our most powerful windows into the physics of the early universe. There exists a well-established theoretical framework connecting the light element abundances to fundamental cosmological parameters; see Refs. Cyburt et al. (2016); Fields et al. (2020); Grohs and Fuller (2023); Navas and others (2024) for recent reviews. The predictions of standard BBN for the primordial abundances of the light elements, particularly helium-4 and deuterium, are in remarkable agreement with measurements, confirming the robustness of the Standard Model of particle physics and cosmology Navas and others (2024). At the same time, because the abundance of these elements depends sensitively on the strength of fundamental interactions and on the cosmic history, BBN offers a unique probe of new physics in the early universe.
Recent years have produced rapid progress in both observational and theoretical aspects of BBN. Measurements of the primordial abundances of helium-4 Aver et al. (2021b); Valerdi et al. (2019); Fernández et al. (2019); Kurichin et al. (2021); Hsyu et al. (2020); Valerdi et al. (2021); Aver et al. (2021a); Navas and others (2024) and deuterium Cooke et al. (2014, 2016); Riemer-Sørensen et al. (2015, 2017); Balashev et al. (2016); Zavarygin et al. (2018); Cooke et al. (2018); Navas and others (2024) have reached unprecedented precision, yielding abundances consistent with Standard Model predictions at the percent level. Experimental efforts, such as those by the LUNA collaboration Mossa and others (2020), have reduced uncertainties in the nuclear reaction rates that govern BBN yields. Moreover, observations of the cosmic microwave background (CMB) now tightly constrain the baryon density Aghanim and others (2020); Louis and others (2025); Camphuis and others (2025), eliminating one of the largest sources of uncertainty in comparing BBN theory with data. Numerical computation of the predictions of BBN for a range of cosmological scenarios can be achieved through various publicly available software packages Arbey (2012); Arbey et al. (2020); Pisanti et al. (2008); Consiglio et al. (2018); Pitrou et al. (2020); Burns et al. (2024); Giovanetti et al. (2025). Together, these advances establish BBN as an incisive tool for testing physics beyond the Standard Model.
@@
Several avenues have been explored in this direction Allahverdi and others (2021). For example, small-scale structure places constraints on modifications to the expansion history at early times Erickcek and Sigurdson (2011); Delos et al. (2018), while gravitational wave backgrounds Caprini and Figueroa (2018) are, in principle, sensitive to the equation of state during very early epochs Watanabe and Komatsu (2006); Barenboim and Park (2016). Studies of the effects of the pre-recombination expansion history using principal component methods (Samsing et al., 2012) and constraints from the matter power spectrum on early dark energy (Sobotka et al., 2024) have provided complementary insights into possible deviations from the standard expansion history. Yet BBN offers a particularly powerful and direct probe, since light element abundance yields are directly sensitive to the expansion rate during the first few minutes after the Big Bang.
Previous studies have shown that BBN abundances can be used to constrain the effective equation of state during nucleosynthesis Carroll and Kaplinghat (2002); Arbey and Mahmoudi (2010); McKeen and Omar (2024), but the steadily improving precision of abundance measurements now makes it possible to go further. Recently, Ref. (An and Gluscevic, 2024) explored a model-independent reconstruction of the expansion and thermal histories constrained by the light element abundances. In this paper, we use the latest abundance data to derive general constraints on the expansion history during BBN. However, unlike Ref. (An and Gluscevic, 2024), we employ principal component analysis to extract the most robust, model-independent features of these constraints, allowing us to place limits on broad classes of early-universe scenarios, including those motivated by early dark energy. In addition, we revisit the long-standing lithium problem, the persistent discrepancy between predicted and observed primordial lithium abundances Fields (2011); Navas and others (2024). While it is well established that uncertainties in nuclear reaction rates cannot resolve this discrepancy Navas and others (2024), we investigate whether a general modification to the expansion history could mitigate the tension.
@@ II Big Bang Nucleosynthesis
As such, BBN plays a pivotal role in constraining both standard cosmological parameters and potential new physics beyond the Standard Model, such as decaying particles, asymmetric reheating, or varying fundamental constants Navas and others (2024). Among the most robust predictions are those for deuterium and helium-4, whose abundances can now be measured with high precision in nearly pristine astrophysical environments. The concordance between BBN predictions and observational data for these elements offers strong support for the standard cosmological model. However, the observed primordial abundance of lithium-7 remains significantly lower than theoretical expectations, a discrepancy known as the “lithium problem,” which may point to underestimated observational errors or hints of new physics Fields (2011); Fields et al. (2020); Navas and others (2024).
@@ II.1 Observational Status
Recent observations of primordial light element abundances continue to support the standard cosmological model for most isotopes while highlighting persistent anomalies. We use the Particle Data Group (PDG) Navas and others (2024) recommended values of observed abundance ratios throughout this work, which are summarized in Table 1. Deuterium remains the most accurately measured primordial isotope, with high-redshift quasar absorption systems yielding D/H = $(2.547\pm 0.029)\times 10^{-5}$, in excellent agreement with BBN predictions when using the CMB-derived baryon density. Similarly, helium-4 determinations from extragalactic HII regions, aided by improved emissivity corrections and infrared line data, support a mass fraction $Y_{p}=0.245\pm 0.003$, matching theoretical expectations. The primordial abundance of helium-3, however, remains poorly constrained due to its detection being limited to metal-rich Galactic environments. Lithium-6, once thought to be overabundant, is now strongly constrained in both Galactic halo stars and the Small Magellanic Cloud (SMC), with no confirmed detections, effectively ruling out a primordial 6Li plateau and supporting the case for stellar depletion.
Lithium-7 remains a significant challenge to the standard BBN paradigm. The Spite plateau in metal-poor halo stars yields a 7Li/H ratio of $(1.6\pm 0.3)\times 10^{-10}$, a factor of three below the BBN+CMB prediction of $(4.72\pm 0.7)\times 10^{-10}$. New observations reveal dispersion and decreasing 7Li levels at metallicities [Fe/H] $\lesssim-2.7$, pointing to stellar depletion as a plausible explanation Fields and Olive (2022). In parallel, stringent upper limits on 6Li/7Li, such as <0.007<0.007 in HD 84937 and <0.1<0.1 in the SMC’s Sk 143 sightline, suggest significant 6Li depletion. Since 6Li is more fragile than 7Li, these findings imply that comparable 7Li destruction is astrophysically feasible, potentially resolving the discrepancy without requiring new physics.
Yet, extragalactic evidence challenges this resolution. The detection of gas-phase 7Li in the SMC with an abundance $A(^{7}\mathrm{Li})\approx 2.2$, unaffected by stellar processes, still falls below BBN expectations Molaro et al. (2024). This implies a primordial shortfall that stellar depletion alone cannot explain. While Galactic 6Li non-detections support an astrophysical solution, the SMC results suggest a possible need for physics beyond the Standard Model. Although stellar depletion offers a compelling explanation within the Milky Way, the lithium problem may ultimately point to deeper cosmological questions.
@@
Table 1: Primordial light element abundances: observations vs. Standard BBN predictions (assuming $\eta_{10}=6.143\pm 0.190$). This work uses the PDG Navas and others (2024) recommended values of observed abundance ratios.
@@ III.1 BBN Calculations
To study the effects of a non-standard expansion history on BBN, we use a modified version of AlterBBN (Arbey, 2012; Arbey et al., 2020). This publicly available code numerically solves the coupled differential equations governing the evolution of light nuclei in the early universe. Boltzmann solvers like AlterBBN are essential tools in cosmology because they compute the time evolution of particle species based on their interactions, decay rates, and the universe’s expansion history. Specifically for BBN, AlterBBN integrates the nuclear reaction network for light elements, such as deuterium, helium-3, helium-4, and lithium-7, while simultaneously solving the Friedmann equation that determines the expansion rate as a function of time $H(t)$.
@@ III.3 Fisher Matrix and Principal Component Analysis
These principal components provide a powerful basis for model-independent constraints: any arbitrary EDE history can be projected onto the eigenbasis, allowing one to reconstruct the constrained and unconstrained features of the model in terms of measurable quantities. The plot of the three eigenmodes are shown in Figure 3. The most strongly constrained mode is degenerate with a uniform increase to the radiation density. However, as seen in the second and third eigenmodes, the primordial abundances are also sensitive to temperature-dependent changes in the expansion history, especially those that affect expansion in the temperature range $10^{9}~\mathrm{K}>T_{\gamma}>3\times 10^{8}$ K. Qualitatively similar conclusions were drawn from the binned expansion history analysis presented in Ref. An and Gluscevic (2024), where the authors also found that modifications to the expansion rate around this temperature range have the most significant impact on light element abundances. The agreement between the PCA-based and binned analyses provides a strong validation that the constraints identified here are robust to the choice of parametrization.
@@ III.4 Markov Chain Monte Carlo Analysis
To explore the full posterior distribution of EDE histories and relevant cosmological parameters, we perform a Markov Chain Monte Carlo (MCMC) analysis using the Cobaya framework Torrado and Lewis (2021). Cobaya (Code for Bayesian Analysis) is a modular, extensible Python-based platform for efficient Bayesian inference, particularly well suited to hierarchical physical models with complex interdependencies.
@@
  • $\eta_{10}\equiv 10^{10}\times\frac{n_{b}}{n_{\gamma}}$: the baryon-to-photon ratio, with a Gaussian prior $\eta_{10}=6.104\pm 0.058$, consistent with the constraint derived from CMB observations Aghanim and others (2020); Navas and others (2024),
  • $\tau_{n}$: the neutron lifetime, with a Gaussian prior $\tau_{n}=878.4\pm 0.5$ s, consistent with laboratory measurements Navas and others (2024),
@@
Each set of parameters defines a complete model of early dark energy. Note that the first principal component of the dark energy density is degenerate with that of the radiation density, typically parameterized by $N_{\mathrm{eff}}$, so we did not separately vary $N_{\mathrm{eff}}$ in our analysis. Several previous studies have shown how observations of primordial abundances can be used to constrain the light relic density Navas and others (2024); Cyburt et al. (2016); Fields et al. (2020); Laguë and Meyers (2020); Dvorkin and others (2022); Yeh et al. (2022).
@@
Sampling is performed using Cobaya’s built-in Metropolis-Hastings algorithm, which automatically groups parameters into fast and slow blocks based on the computational cost of the components that depend on them Lewis and Bridle (2002); Lewis (2013); Neal (2005); Torrado and Lewis (2021). This automatic parameter blocking and caching of intermediate results greatly improves sampling efficiency. Chain convergence is monitored using the Gelman-Rubin R1R-1 statistic, with a typical stopping criterion of R1<0.01R-1<0.01.
Posterior samples are processed using GetDist Lewis (2025), which interfaces seamlessly with Cobaya’s output, providing marginal distributions, credible intervals, and covariance matrices. In particular, we analyze the posterior distributions of the eigenmode amplitudes $\alpha_{i}$, and project the results back into temperature space to reconstruct allowed EDE histories and identify epochs most sensitive to BBN constraints. We use the observational constraints on primordial abundances from the PDG Navas and others (2024), summarized above in Table 1.
@@ IV Results
Figure 4 shows the allowed EDE histories consistent with the observed primordial helium-4 and deuterium abundances. In this scenario, the EDE component is generally constrained to remain subdominant throughout the epoch of BBN, except in the narrow temperature range $3\times 10^{8}\,\mathrm{K}\lesssim T\lesssim 5\times 10^{8}\,\mathrm{K}$, where constraints are weaker, and an EDE component with energy density comparable to the radiation density is permitted. A second poorly constrained region is also visible around or just after the deuterium bottleneck, indicating a localized allowance for additional EDE energy density during that stage. It should be noted, however, that a localized spike in energy density is unlikely to be realized in a self-consistent physical model of EDE. For example, the equation of state of any fluid obeying the null energy condition satisfies $w\geq-1$, and thus has an energy density that either remains constant or dilutes with the expansion of space. Furthermore, a free scalar field has $w=1$ with an energy density scaling as $\rho\propto a^{-6}$, while the presence of a non-vanishing potential would lead to a slower decrease in energy density. Thus the bounds we present here should be interpreted as indicating the regions where observational constraints are more or less constraining, while bearing in mind that it may not be possible to saturate the upper limits in the less well-constrained regions in a self-consistent model.
These results are broadly consistent with the binned reconstruction of the expansion history presented in Ref. An and Gluscevic (2024), confirming that the light-element abundances primarily constrain deviations in the Hubble rate near the onset of nucleosynthesis. The PCA framework developed here extends that earlier analysis by providing a continuous, model-independent mapping of the allowed expansion histories, revealing not only where the expansion rate is tightly constrained but also which temperature ranges remain weakly bounded by current data.
@@
The analysis presented here also allows us to determine how future data might impact our ability to constrain the early expansion history. Figure 5 presents the corresponding histories obtained after including a forecasted $1\%$ constraint on the helium-3 abundance. The addition of this prior produces minimal qualitative change in the allowed EDE behavior. As before, EDE is required to remain largely subdominant, with the same narrow windows that are more weakly constrained between $3\times 10^{8}$ K and $5\times 10^{8}$ K and near or following the deuterium bottleneck.
In contrast, Figure 6 illustrates the allowed EDE histories when only the helium mass fraction and deuterium abundance are constrained. At the same time, the prior on $\eta_{10}$ is relaxed to a flat range $[1,20]$. In this case, constraints are much weaker throughout the entire BBN epoch, allowing energy densities several orders of magnitude higher than in the previous scenarios.
However, this freedom is accompanied by extreme values of $\eta_{10}$ that are in strong tension with the precise constraints from CMB observations, highlighting the importance of an informed prior on the baryon-to-photon ratio.